Earley Corporation issued perpetual preferred stock with an 11% annual dividend. The stock currently yields 7%, and its par value is $100. Round your answers to the nearest cent. What is the stock's value? $

Suppose interest rates rise and pull the preferred stock's yield up to 11%. What is its new market value? $
Answer #1

Annual dividend=100*11%=11

Stock value=Annual dividend/Current yield

a.Stock value=11/0.07=$157.14(Approx).

b.Stock value=11/0.11=$100